Ed is a runner and he runs a 8 km loop every day.  The first 4 km, he
runs at 12km/hr.  He runs much slower on the way home.  If it takes him 1 hour in total to
run the loop, how fast is he running for the last 4 km?

Answer by jim_thompson5910(35256)

The formula we'll use here is
d = r*t
where,
d = distance
r = rate (aka speed)
t = time

If Ed runs 4 km and his speed is 12 km/hr, then d = 4 and r = 12. Plug these values into the formula above and solve for t
d = r*t
4 = 12*t
4/12 = 12*t/12
1/3 = t
t = 1/3

So it takes Ed 1/3 of an hour, which is (1/3)*60 = 20 minutes to run the 4 km.

20 minutes pass by when Ed runs the first 4 km, leaving 60 - 20 = 40 minutes left over.

40 minutes = 40*(1/60) = 40/60 = 4/6 = 2/3 of an hour

So we'll use t = 2/3 and d = 4 to find the rate (r)

d = r*t
4 = r*(2/3)
3*4 = 3*r*(2/3) ... multiply both sides by 3
12 = 2r
2r = 12
2r/2 = 12/2 ... divide both sides by 2
r = 6

Therefore, Ed's speed on the last four kilometers is 6 km/hr which is the final answer.